Present Value of Bonds Payable; Premium Mason Co. issued $610,000 of five-year, 12% bonds with interest payable semiannually, at a market (effective) interest rate of 10%. Determine the present value of the bonds payable, using the present value tables in Exhibit 4 and Exhibit 5. Round to the nearest dollar.
Answer #1

DETERMINE PRESENT VALUE OF BONDS PAYABLE :

interest rate = 12/2 =6%

market rate = 10/2 = 5%

Time = 5*2 = 10 years

Interest = 610000*6% = 36600

Present value of bonds payable = present value of interest +present value of maturity

                                             = (36600* 7.7217) +(610000*0.61391)

Present value of bonds payable = $657100